# Integrating New Technologies into Science:

### 15 Dec 2023 v1

@@ 4.2.2 AI re-use
The results on the role of computer scientists are somewhat peculiar. Regarding the first AI paper, the presence of computer scientists in the prior collaboration network (CS col.) is positively associated with AI reuse. But the presence of computer scientists as co-authors in the first AI paper (CS aut.) plays exactly the opposite effect. [^14] In order to ensure that this result does not arise from ignoring non-linear effects of individual variables or their interaction, we tried various alternative specifications with and without squared terms and with varying sets of variables. This result appeared to be robust, and even stronger when adding squared terms. A domain scientist collaborating with, say, three computer scientists in the first AI paper, compared to no computer scientists in the team, has an approximately 10% lower odds-ratio for AI reuse. Although the estimated association is of relatively small magnitude compared to other factors, a brief discussion seems warranted. We propose the following explanation. Collaboration with computer scientists is likely to be instrumental in the diffusion of AI, as multi-disciplinary teams seem well suited to apply and adapt AI tools from computer sciences to research problems in other domains. Our data, indeed, support the relevance of collaboration between computer scientists and researchers of other domains in AI adoption. It is also clear, however, that such interdisciplinary teams may face various hurdles in their formation and subsequent life-cycle (see, e.g., Fiore, 2008). For instance, in their analysis of scientific collaboration among Stanford faculty members, Dahlander and McFarland (2013) find that intellectual dissimilarity has a negative effect not only on first-tie formation but also on collaboration persistence. Similarly, our findings show that collaboration between computer scientists and domain scientists that led to the first AI paper is in general difficult to maintain and is unlikely to lead to persistent AI adoption by domain scientists.

### 22 Feb 2024 v2

@@ 4.2.2 AI re-use
The results on the role of computer scientists are somewhat peculiar. Regarding the first AI paper, the presence of computer scientists in the prior collaboration network (CS col.) is positively associated with AI reuse. But the presence of computer scientists as co-authors in the first AI paper (CS aut.) plays exactly the opposite effect. A domain scientist collaborating with, say, three computer scientists in the first AI paper, compared to no computer scientists in the team, has an approximately 10% lower odds-ratio for AI reuse. Although the estimated association is of relatively small magnitude compared to other factors, a brief discussion seems warranted. We propose the following explanation. Collaboration with computer scientists is likely to be instrumental in the diffusion of AI, as multi-disciplinary teams seem well suited to apply and adapt AI tools from computer sciences to research problems in other domains. Our data, indeed, support the relevance of collaboration between computer scientists and researchers of other domains in AI adoption. It is also clear, however, that such interdisciplinary teams may face various hurdles in their formation and subsequent life-cycle (see, e.g., Fiore, 2008). For instance, in their analysis of scientific collaboration among Stanford faculty members, Dahlander and McFarland (2013) find that intellectual dissimilarity has a negative effect not only on first-tie formation but also on collaboration persistence. Similarly, our findings show that collaboration between computer scientists and domain scientists that led to the first AI paper is in general difficult to maintain and is unlikely to lead to persistent AI adoption by domain scientists.